In an AP, the sum of the first three terms is 18, and the sum of the squares of these terms is 126. Find the three terms.

Answer by richwmiller(17219)

a+b+c=18,
a^2+b^2+c^2=126,
b-a=d
c-b=d
a = 3, b = 6, c = 9, d = 3
or
a = 9, b = 6, c = 3, d = -3